[20pt] 8. Find the solution of the IVP r = 14 + x2 + ta?, x(0) = 0.
Answer #1

(8) Solution: Given IVP is x=1+t+x² + tx,x(0) = 0 dx = (1+t)+(1+t)x 2 dt dx dt = (1+t)(1+x2) Taking variable and seperable,w